i)                For a system of 3 distinguishable particles distributed in two similar compartments, the total number of microstates corresponding to the condition that no compartment should remain empty are______.

a)     4, b) 6, c) 3, d) 9.



1
Expert's answer
2021-04-26T17:08:55-0400

As particles are distinguishable, the possible states depend on the position of each particle, where the position means just in which compartment they are situated. The number of possible states is "2^3=8", as it is "2 \text{ states of particle A} \times 2 \text{ states of particle B}\times 2 \text{ states of particle C}". Now we need to exclude the states where one compartiment is empty : there is exactly two such states :

  • all particles are in the first compartment
  • all particles are in the second compartment

Therefore there is "8-2=6" microstates where no compartment is empty.

The answer is b.